OpenStudy (anonymous):

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? 4x - y = 13 4x + y = 13 4x + y = -13

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

HI? can u help?

OpenStudy (misty1212):

you answer has to have \(4x+y\) in it right?

OpenStudy (misty1212):

not \(4x-y\) so you can forget the first one

OpenStudy (anonymous):

idek

OpenStudy (misty1212):

the line you are given in the question is \[\huge 4x+y-2=0\] which has \(4x+y\) in it

OpenStudy (anonymous):

okay

OpenStudy (misty1212):

and you want it to pass through \((4,-3)\) right?

OpenStudy (anonymous):

Yes.

OpenStudy (misty1212):

the \(4\) is the \(x\) and the \(-3\) is the \(y\) so lets substitute \(4\) for \(x\) and \(-3\) for \(y\) in \[4x+y\] and see what we get

OpenStudy (anonymous):

I got 11=0

OpenStudy (misty1212):

i.e. compute \[4\times 4+(-3)\]

OpenStudy (misty1212):

\(11=0\)??!!

OpenStudy (anonymous):

I'm literally confused.

OpenStudy (misty1212):

what is \[\huge 4\times 4-3\]?

OpenStudy (anonymous):

13

OpenStudy (misty1212):

yeah if you think \(11=0\) you must be confuffled

OpenStudy (misty1212):

right \(13\)

OpenStudy (misty1212):

so it must be \[4x+y=13\]

OpenStudy (anonymous):

TYSM!!!!!

OpenStudy (triciaal):

parallel lines have the same slope 4 x + y - 2 = 0 slope is -4 passing through (4, -3) y --3 = -4(x - 4) y + 3 = -4x + 16 standard form 4 x + y = 13